silvernight269:

A segment is divided into two parts having lengths in a ration of 7:2. Find the lengths of each part, if the difference between the segment lengths is 15 inches.

silvernight269:

@Shadow

SmokeyBrown:

Hey, let's take the lengths of these two segments, call them x and y. Say x is the bigger of the two. So, we know that x = 7.2y We also know that x-y = 15 Now we have a system of equations. If you combine them, you can solve for y, then use that to solve for x

silvernight269:

no its 7x and 2x cus if the ratio is 7:2

SmokeyBrown:

Oh, yes, I misread 7:2 as 7.2 I must be getting old, rip But yeah, in that case the first equation would be 2x = 7y

silvernight269:

ok....

SmokeyBrown:

Yup, so then you can simplify that to x = 3.5y Then we take our other equation, x - y = 15. We can rewrite it as x = y + 15. And now we have two equations that tell us x is equal to something. Well, if a=b and a=c, then b=c. So, we can combine these two equations like 3.5y = y + 15 Now we can isolate y 2.5y = 15 y = 15/2.5 = 6 Now, since y = 6, and x - y = 15, x must be 21 Does that match up? Well, 21 * 2 is 42. And 6 * 7 is 42. So, our equations check out.

silvernight269:

i dont think you can simplify 7y=2x to x=3.5y cus of the two different variables

silvernight269:

@SmokeyBrown

SmokeyBrown:

Oh, nah, it's cool. The variables aren't actually interacting with each other, like if we tried to add or subtract them. We're just dividing the coefficients (the numbers multiplying the variables)

silvernight269:

ok

silvernight269:

but the difference of both length is 15

SmokeyBrown:

Yeah, that's true. So, x - y = 15, if x is larger. And the results we got match, since 21 - 6 does give a difference of 15

silvernight269:

ok